Toxaway Company is a merchandiser that segments its business into two divisions-Commercial and Residential. The companys accReg 1 Reg 2A Reg 2B Reg 3 Reg 4 Reg 5 Reg 6 Req 7 Based on a review of the interns segmented income statement, how much of tReg 1 Req 2A Reg 2B Req3 Reg 4 Reg 5 Reg 6 Req 7 Based on a review of the interns segmented income statement, which of the f| Regi Req2A Reg 1 Reg 2A Req 2B Reqze Req3 Reg 3 Rea> Reg 4 Rega Reg 5 Reas ReaG Reg 6 Reg 7 Do you agree with the internsReq 1 Reg 2A Req 2B Req3 | Reg 4 Req 5 Redo the interns segmented income statement using the contribution format. Toxaway CoCompute the companywide break-even point in dollar sales. (Round intermediate calculations to 3 decimal places and final answReg 1 Req 1 Reg 2A Reg za Reg 2B Reg 2 Reg3 Reg 4 Reg 4. Reg 5 Reg5 Rego Reg 6 Reg 7 Req 7 Compute the break-even point in doReg 1 Req 2A Req 2B Req3 Reg 4 Reg 5 Reg 6 Req 7 Assume the company decided to pay its sales representatives in the Commercia
Answer #1
1 No.
Absorption format is used to report to the outside parties.
It is appropriate to use contribution format for the decision making purpose
2
a. Commercial Residential
Total selling and administrative expenses 132000 164000
Less:
Sales commission (Sales*10%) 32000 64000
Specific fixed expenses 78000 56000
110000 120000
Common fixed expenses allocated 22000 44000
b. Ratio of common fixed expenses allocated=22000:44000 or 1:2
Ratio of sales=320000:640000 or 1:2
Hence, sales would have been used as the allocation base
3 No.
Common fixed expenses should not be allpcated to products since it is not directly related to the product.
It will distort the operating income
4 Commercial Residential Total company
Sales A 320000 640000 960000
Less:Variable expenses
Cost of goods sold 169600 467200 636800
Selling and administrative expenses (10% of sales) 32000 64000 96000
Total variable expenses B 201600 531200 732800
Contribution margin C=A-B 118400 108800 227200
Contribution margin ratio D=C/A 0.37 0.17 0.24
Less:Fixed selling and administrative expenses E 78000 56000 134000
Segment operating income F=C-E 40400 52800 93200
5 Companywide break-even point in dollar sales=Total fixed expenses including common fixed expenses/Total contribution margin ratio=(134000+66000)/0.24=833333.3=$ 833333
6 Break-even point for commercial division=Specific fixed expenses/Contribution margin ratio=78000/0.37=$ 210811
Break-even point for residential division=Specific fixed expenses/Contribution margin ratio=56000/0.17=$ 329412
7
Commercial Residential Total company
Sales A 320000 640000 960000
Less:Variable expenses
Cost of goods sold 169600 467200 636800
Selling and administrative expenses (5% of sales) 16000 32000 48000
Total variable expenses B 185600 499200 684800
Contribution margin C=A-B 134400 140800 275200
Contribution margin ratio D=C/A 0.42 0.22 0.29
Less:Fixed selling and administrative expenses E 103000 106000 209000
(78000+25000) (56000+50000)
Segment operating income F=C-E 31400 34800 66200
Break-even point for commercial division=Specific fixed expenses/Contribution margin ratio=103000/0.42=$ 245238
Break-even point for residential division=Specific fixed expenses/Contribution margin ratio=106000/0.22=$ 481818